Circular No. 7223

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA AQUILAE 1999 S. Nakano, Sumoto, Japan, reports the discovery by Akihiko Tago,

Tsuyama, of an apparent nova (mag 8.8) on two TMax-400 films taken on July 13.558, 13.560, 13.565, and 13.567 UT with a 55-mm f/3 camera lens. Tago places the object at α = 19^{h}07^{m}40^{s}, δ = +12^{o}32′.2 (equinox 2000.0), and he notes that nothing is visible at this location on his films of July 5 and 9 (limiting mag 11 and 10.5, respectively). On a film taken on July 14.5, Tago derived mag ∼ 8.9 for the star. Nakano also reports the following position from a CCD image obtained by A. Nakamura, Kuma Kogen, on July 14.513: α = 19^{h}07^{m}36^{s}.90, δ = +12^{o}31′26′′.2. The closest star on the Digital Sky Survey is one of mag ∼ 18 located ∼ 4′′ northeast of Nakamura’s position; a red star (R = 16.4) located 8′′ southwest of Nakamura’s position is listed in the USNO A1.0 catalogue.

K. Ayani and T. Kawabata, Bisei Astronomical Observatory (BAO), Okayama, report: “We obtained a low-resolution spectrum (range 470– 680 nm) of Tago’s suspected nova on July 14.6 UT with the BAO 1.01-m telescope. Strong, broad Hα and Hβ lines are visible. The Hα line has FWHM ∼ 3400 km/s and equivalent width ∼ 50 nm.”

THE EDGAR WILSON AWARD With reference to IAUC 6936, the Smithsonian Astrophysical Observa-

tory wishes to announce that the 1999 Edgar Wilson Award for the discovery of comets by amateur astronomers is being divided among the following six individuals or groups: Peter Williams, Heathcote, N.S.W., Australia, for C/1998 P1; Roy A. Tucker, Tucson, AZ, U.S.A., for P/1998 QP_{54}; Michael Jäger, Weissenkirchen i.d. Wachau, Austria, for P/1998 U3; Justin Tilbrook, Clare, S.A., Australia, for C/1999 A1; Korado Korlević and Mario Jurić, Višnjan, Croatia, for P/1999 DN_{3}; and Steven Lee, Coonabarabran, N.S.W., Australia, for C/1999 H1.

SUPERNOVA 1999by IN NGC 2841 CCD magnitudes by D. Hanžl, Brno, Czech Republic (comparison star

GSC 3431.625, using Tycho magnitudes): June 2.868 UT, V = 14.70 ± 0.03, V − R = +0.30 ± 0.05, V − I = +0.8 ± 0.1; July 4.852, V = 16.1 ± 0.1, V − R = +0.21 ± 0.05; 5.860, R = 15.8 ± 0.2; 5.871, I = 15.3 ± 0.1.

1999 July 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green